Find the following probabilities:

(a) P(z > -0.78)

(b) P(0.87 < z < 2.35)
Answer #1

solution

(A)P(z > -0.78) =1 - P(z <-0.78 )

Using z table,
= 1 -0.2177

= 0.7823   

probability =0.7823

(B) P(0.87 < z < 2.35)

= P(Z <2.35 ) - P(Z <0.87 )

Using z table,  

= 0.9906-0.8078   

probability=0.1828